A mother figures that her son will forget his homework about 20% of the time. What is the probability that he will forget his homework at least once in the next 10 school days?


1
Expert's answer
2021-03-30T07:49:00-0400

This is a binomial probability with n=10, p=0.2.

"P(X<=1)=1-P(X=0)=1-0.8^{10}=0.8926."